OpenStudy (anonymous):

Please assist. Thank you The tensile strength (the maximum tensile stress it can support without breaking) for a certain steel wire is 3000 MN/m2. What is the maximum load that can be applied to a wire with a diameter of 3.0 mm made of this steel without breaking the wire?

OpenStudy (aaronq):

It seems that you would need to find the area of the cross-section of the wire and multiply it by the tensile strength. \(\sf A_{circle}=\pi r^2=\pi (\dfrac{d}{2})^2\) Max load =\(\sf 3000~MN/m^2*\pi (\dfrac{d}{2})^2\)